What is the lowest amount of work required to pull a 950 kilogram granite block up an inclined plane tilted at 9.0° above horizontal if the inclined plane is 710 meters long?

between 8.0 x 105 J and 8.5 x 105 J
between 8.5 x 105 J and 9.0 x 105 J
between 9.0 x 105 J and 9.5 x 105 J
greater than 9.5 x 105 J
smaller than 7.0 x 105 J
between 7.0 x 105 J and 7.5 x 105 J
between 7.5 x 105 J and 8.0 x 105 J
Answer #1

here,

the mass of block , m = 950 kg

theta = 9 degree

the length of incline , l = 710 m

the minimum work done required , W = m * g * l * sin(theta)

W = 950 * 9.81 * 710 * sin(9) J

W = 10.4 * 10^5 J

correct option is (greater than 9.5 * 10^5 J)